OpenStudy (aimmebgood):

Factor completely x^(2)(x+4)(-6x(x+4)+9(x+4). Ive gotten to (x+4)(X^(2) - 6X +9). I dont know where to go from here.

OpenStudy (anonymous):

recognize the second factor as a perfect square

OpenStudy (anonymous):

\[factor: x^{2}-6x+9\]

OpenStudy (anonymous):

the factors are 0,-4,2/3

OpenStudy (anonymous):

i think for the factors of \[f(x)=a \times b \times c\] a=0,b=0,c=0 separately where a,b,c are the functions of x

OpenStudy (aimmebgood):

that doesnt really help.

OpenStudy (anonymous):

maybe this will help \[(a-b)^2=a^2-2ab+b^2\]

OpenStudy (anonymous):

replace \(a\) by \(x\) and \(b\) by \(3\)

OpenStudy (anonymous):

making \[x^2-6x+9=(x-3)^2\]

OpenStudy (anonymous):

i mean to say assume \[a=x ^{2},b=x+4,c=(-6x(x+4)+9(x+4)\] and then find the zeroes

OpenStudy (anonymous):

and a "final answer" of \[(x+4)(x-3)^2\] to the problem

OpenStudy (anonymous):

it says nothing about finding zeros it says to factor

OpenStudy (anonymous):

@aimmeBgood your first step was good next step is to rewrite \(x^2-6x+9\) as \((x-3)^2\) hope it is clear

OpenStudy (anonymous):

the factors thus could be found as x+4 x-2/3 sorry but i interpret the question as \[x ^{2}(x+4)(-6x(x+4)+9(x+4))\]

OpenStudy (anonymous):

sorry its x+3/2 instead of x-2/3

OpenStudy (anonymous):

also i forgot to mention that x is also its factor